Write the equation of the line with slope 5/3 which passes through the point (3, -2).
Answer by mananth(16946)

m= 5/ 3

Plug value of the slope and point ( 3 , -2 ) in
Y = m x + b
-2.00 = 5 + b
b= -2 - 5
b= -7
So the equation will be
Y = 5/3 x -7
m.ananth@hotmail.ca